# Resolving Quantum Criticality in the Honeycomb Hubbard Model

### v1

@@ IV.1 Projector determinant quantum Monte Carlo
Systematic errors in PQMC arise primarily from the Trotter discretization $\Delta_{\tau}$ and the finite projection length $\beta$. We control these errors by performing convergence tests for both models (see Supplemental Material [47] Section S4 for details). For the Hubbard model, we adopt $\Delta_{\tau}t=0.1$ [19, 18, 20], whereas for the $t$-$V$ model, a smaller time step $\Delta_{\tau}t=0.05$ is required [15]. The projection length is scaled with the linear system size as $\beta t=L+12$ for both models, which is sufficient to render systematic errors smaller than statistical fluctuations.
@@ IV.2 Submatrix update algorithm for PQMC
We first briefly review the traditional fast update algorithm for $T\equiv(LR)^{-1}$. More details can be found in Supplemental Material [47] Section S1. After tracing out the fermionic degrees of freedom, the Monte Carlo weight for configuration $\mathbf{s}$ is proportional to a fermionic determinant $\det[L(\tau)R(\tau)]$, where $R(\tau)$ (size $N\times N_{p}$) and $L(\tau)$ (size $N_{p}\times N$) are rectangular matrices encoding the imaginary-time propagation from the trial state to time slice $\tau$ from 0 and $\beta$, respectively. Here $N$ typically denotes the total number of lattice sites and $N_{p}$ is the number of particles per spin block.
@@
where $e_{x_{j}}=[0,\ldots,0,1,0,\ldots,0]^{T}$ is the $x_{j}$-th unit vector of length $N$ (several properties of the index matrices are discussed in Supplemental Material [47] Section S1). The acceptance ratio is given by $r=\det[I_{k}+VU]$. The per-move cost is $\mathcal{O}(N_{p}^{2})$ and the per-time-slice cost is $\mathcal{O}(N_{p}^{2}N)$.
While this fast update formula enables efficient per-move computation, it relies heavily on rank-$k$ matrix-vector operations (BLAS-1&2) that are not cache-friendly. To overcome this bottleneck, delayed update algorithms, accumulating multiple local updates before performing a single, cache-efficient, higher-rank batch operation (BLAS-3), were proposed [44, 45, 46]. In Supplemental Material [47] Section S2, we provide a self-contained review of delay-G [44], submatrix-G [45], and delay-T [46] algorithms.
@@
The most essential insight for designing submatrix-T update is that $L$, $R$ and $T$ can find their general term formulas after arbitrary number of accepted local updates. Within our sweep strategy, at a given time slice, $L$ does not update so $L^{(i)}=L^{(0)}=L$ (c.f. Eq. (S14) in Supplemental Material [47] Section S1). Since different local updates operate on disjoint lattice sites, the matrix $R$ can be updated by successive multiplication:
@@
where we further factorize the $k\times k$ determinant with matrices $\mathcal{V}_{k\times k}^{(i+1)}\equiv P_{k\times N}^{(i+1)}F^{(i)}P_{N\times k}^{(i+1)}$ and $D_{k\times k}^{(i+1)}\equiv P_{k\times N}^{(i+1)}\Delta^{(i+1)}P_{N\times k}^{(i+1)}$. The $N\times N$ matrix $F^{(i)}\equiv R^{(i)}T^{(i)}L$ has a similar role as the Green’s function $G^{(i)}$ (which equals to $(I-F^{(i)})$), and it is hired as the update object in the submatrix-G algorithm (see Supplemental Material [47] Section S2 for detailed formulations). Its general term formula follows from Eqs. (12) and (14):
@@
Figure 6 benchmarks the update time of submatrix-T algorithm against fast update and other existing delayed update algorithms. Submatrix-T consistently outperforms all alternatives. For the $L=36$ honeycomb lattice in single-threaded execution, submatrix-T achieves more than $24\times$ speedup over fast update, and this speedup factor increases with system size, reaching higher acceleration for larger lattices. The superior performance stems from replacing cache-inefficient BLAS-1&2 operations with batched BLAS-3 routines, which not only improves cache utilization but also provides better amenability to parallel acceleration on multi-threaded (e.g., OpenMP) and GPU architectures. For implementation details including the workflow flowchart, computational complexity analysis, hardware specifications, and comprehensive $n_{d}$ optimization results, see Supplemental Material [47] Section S3. As a low-level optimization, our submatrix-T algorithm can be broadly applied to accelerate fermionic computations within the PQMC framework, including entanglement measures [53, 54, 55, 56], stabilizer Rényi entropy [57], and finite-temperature schemes evaluating Fock-state expectations [58, 59].
@@ IV.3 Data collapse analysis
To avoid choosing an explicit ansatz for $F$, we adopt the Bayesian scaling analysis (BSA) based on Gaussian-process regression [60]. The scaling function $F$ is modeled as a Gaussian process defined by its kernel hyperparameters. The physical scaling parameters ($X_{c}$, $c_{1}$, $c_{2}$) and the hyperparameters are fitted jointly; in what follows, we refer to all of them collectively as the fitting parameters. We use the reduced chi-squared $\chi^{2}_{\mathrm{red}}=\chi^{2}/(N_{\mathrm{data}}-N_{\mathrm{free}})$ to evaluate the goodness of fit, where $N_{\mathrm{data}}$ is the number of data points in the fitting window and $N_{\mathrm{free}}$ is the number of unfixed fitting parameters. The chi-squared is computed as $\chi^{2}=\sum_{i}\left[F(X_{i})-Y_{i}\right]^{2}/E_{i}^{2}$, where $E_{i}^{2}=\sigma_{Y_{i}}^{2}+\left[F^{\prime}(X_{i})\right]^{2}\sigma_{X_{i}}^{2}-2F^{\prime}(X_{i})\sigma_{X_{i}Y_{i}}$ is the effective variance that accounts for uncertainties in both $X$ and $Y$ data, as well as their covariance $\sigma_{X_{i}Y_{i}}$ [61, 62]. We provide more details in the $\chi^{2}_{\mathrm{red}}$ calculation of the three kinds of collapses in Supplemental Material [47] Section S9.
@@ IV.4 Acknowledgments
We thank J. A. Gracey and T. C. Lang for helpful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the National Natural Science Foundation of China (Grants No. 12447103, No. 12274289 and No. 125B2077), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and Shanghai Jiao Tong University 2030 Initiative. F.-H. W. is supported by T.D. Lee scholarship. The computations in this paper were partially run on the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University.

### v2

@@ IV.1 Projector determinant quantum Monte Carlo
Systematic errors in PQMC arise primarily from the Trotter discretization $\Delta_{\tau}$ and the finite projection length $\beta$. We control these errors by performing convergence tests for both models (see Supplementary Note 4 for details). For the Hubbard model, we adopt $\Delta_{\tau}t=0.1$ [22, 21, 23], whereas for the $t$-$V$ model, a smaller time step $\Delta_{\tau}t=0.05$ is required [18]. The projection length is scaled with the linear system size as $\beta t=L+12$ for both models, which is sufficient to render systematic errors smaller than statistical fluctuations.
@@ IV.2 Submatrix update algorithm for PQMC
We first briefly review the traditional fast update algorithm for $T\equiv(LR)^{-1}$. More details can be found in Supplementary Note 1. After tracing out the fermionic degrees of freedom, the Monte Carlo weight for configuration $\mathbf{s}$ is proportional to a fermionic determinant $\det[L(\tau)R(\tau)]$, where $R(\tau)$ (size $N\times N_{p}$) and $L(\tau)$ (size $N_{p}\times N$) are rectangular matrices encoding the imaginary-time propagation from the trial state to time slice $\tau$ from 0 and $\beta$, respectively. Here $N$ typically denotes the total number of lattice sites and $N_{p}$ is the number of particles per spin block.
@@
where $e_{x_{j}}=[0,\ldots,0,1,0,\ldots,0]^{T}$ is the $x_{j}$-th unit vector of length $N$ (several properties of the index matrices are discussed in Supplementary Note 1). The acceptance ratio is given by $r=\det[I_{k}+VU]$. The per-move cost is $\mathcal{O}(N_{p}^{2})$ and the per-time-slice cost is $\mathcal{O}(N_{p}^{2}N)$.
While this fast update formula enables efficient per-move computation, it relies heavily on rank-$k$ matrix-vector operations (BLAS-1&2) that are not cache-friendly. To overcome this bottleneck, delayed update algorithms, accumulating multiple local updates before performing a single, cache-efficient, higher-rank batch operation (BLAS-3), were proposed [54, 55, 56]. In Supplementary Note 2, we provide a self-contained review of delay-G [54], submatrix-G [55], and delay-T [56] algorithms.
@@
The most essential insight for designing submatrix-T update is that $L$, $R$ and $T$ can find their general term formulas after arbitrary number of accepted local updates. Within our sweep strategy, at a given time slice, $L$ does not update so $L^{(i)}=L^{(0)}=L$ (c.f. Eq. (S14) in Supplementary Note 1). Since different local updates operate on disjoint lattice sites, the matrix $R$ can be updated by successive multiplication:
@@
where we further factorize the $k\times k$ determinant with matrices $\mathcal{V}_{k\times k}^{(i+1)}\equiv P_{k\times N}^{(i+1)}F^{(i)}P_{N\times k}^{(i+1)}$ and $D_{k\times k}^{(i+1)}\equiv P_{k\times N}^{(i+1)}\Delta^{(i+1)}P_{N\times k}^{(i+1)}$. The $N\times N$ matrix $F^{(i)}\equiv R^{(i)}T^{(i)}L$ has a similar role as the Green’s function $G^{(i)}$ (which equals to $(I-F^{(i)})$), and it is hired as the update object in the submatrix-G algorithm (see Supplementary Note 2 for detailed formulations). Its general term formula follows from Eqs. (12) and (14):
@@
Figure 6 benchmarks the update time of submatrix-T algorithm against fast update and other existing delayed update algorithms. Submatrix-T consistently outperforms all alternatives. For the $L=36$ honeycomb lattice in single-threaded execution, submatrix-T achieves more than $24\times$ speedup over fast update, and this speedup factor increases with system size, reaching higher acceleration for larger lattices. The superior performance stems from replacing cache-inefficient BLAS-1&2 operations with batched BLAS-3 routines, which not only improves cache utilization but also provides better amenability to parallel acceleration on multi-threaded (e.g., OpenMP) and GPU architectures. For implementation details including the workflow flowchart, computational complexity analysis, hardware specifications, and comprehensive $n_{d}$ optimization results, see Supplementary Note 3. As a low-level optimization, our submatrix-T algorithm can be broadly applied to accelerate fermionic computations within the PQMC framework, including entanglement measures [57, 58, 59, 60], stabilizer Rényi entropy [61], and finite-temperature schemes evaluating Fock-state expectations [62, 63].
@@ IV.3 Data collapse analysis
To avoid choosing an explicit ansatz for $F$, we adopt the Bayesian scaling analysis (BSA) based on Gaussian-process regression [64]. The scaling function $F$ is modeled as a Gaussian process defined by its kernel hyperparameters. The physical scaling parameters ($X_{c}$, $c_{1}$, $c_{2}$) and the hyperparameters are fitted jointly; in what follows, we refer to all of them collectively as the fitting parameters. We use the reduced chi-squared $\chi^{2}_{\mathrm{red}}=\chi^{2}/(N_{\mathrm{data}}-N_{\mathrm{free}})$ to evaluate the goodness of fit, where $N_{\mathrm{data}}$ is the number of data points in the fitting window and $N_{\mathrm{free}}$ is the number of unfixed fitting parameters. The chi-squared is computed as $\chi^{2}=\sum_{i}\left[F(X_{i})-Y_{i}\right]^{2}/E_{i}^{2}$, where $E_{i}^{2}=\sigma_{Y_{i}}^{2}+\left[F^{\prime}(X_{i})\right]^{2}\sigma_{X_{i}}^{2}-2F^{\prime}(X_{i})\sigma_{X_{i}Y_{i}}$ is the effective variance that accounts for uncertainties in both $X$ and $Y$ data, as well as their covariance $\sigma_{X_{i}Y_{i}}$ [65, 66]. We provide more details in the $\chi^{2}_{\mathrm{red}}$ calculation of the three kinds of collapses in Supplementary Note 8.
@@ IV.4 Acknowledgments
We thank J. A. Gracey, T. C. Lang and M. Scherer for helpful discussions. This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the National Natural Science Foundation of China (Grants No. 12447103, No. 12274289 and No. 125B2077), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and Shanghai Jiao Tong University 2030 Initiative. F.-H. W. is supported by T.D. Lee scholarship. The computations in this paper were partially run on the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University.